John and dick had a collection of 79 books. After John gave 15 books dick, he had 5 books fewer than dick. How many books did John have at first ?
1
Expert's answer
2012-09-04T10:03:59-0400
Let John had X books and Dick had Y books.
X + Y = 79
(X - 15) + 5 = (Y +
15)

Y = 79 - X
X - 10 = 79 - X + 15

2*X = 79 + 15 + 10
2*X =
104
X = 52

Answer: 52.